Need help converting word problem into algebraic expresion One-half of n diminished by six equals one wrote it as 1/2(n)-6=1 but unable to get correct
Answer by checkley77(12844)

N/2-6=11+6
N/2=7 CROSS MULTIPLY.
N=2*7
N=14
PROOF:
14/2-6=1
7-6=1
1=1